Foxie has a budget of P2,000 monthly which she spends on lunch @ P80.00 and 2 snacks equivalent to P50.00 daily. What is the combination of lunch and snacks will she have to maximize satisfaction? If she will spend all her income on lunch, how many days will her budget last? How many daily snacks can she enjoy? Determine the total no. of snacks she can have for her budget. Show your solution. Illustrate the indifference curve and budget lines for the two expenditures.



1
Expert's answer
2022-02-10T15:33:14-0500

"\frac{2000}{80}=25"  days will spend all her income on lunch


"\frac{2000}{50}=40" daily snacks can she enjoy


80x+50y=2000